40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 1 19:26:09 2012, jamie, Update, Alignment, aligned, this time without the crying 
    Reply  Thu Nov 1 19:28:48 2012, Jenne, Update, Alignment, aligned, AS beam on camera 
Message ID: 7657     Entry time: Thu Nov 1 19:26:09 2012     Reply to this: 7658
Author: jamie 
Type: Update 
Category: Alignment 
Subject: aligned, this time without the crying 

Jamie, Jenne, Nic, Manasa, Raji, Ayaka, Den

We basically walked through the entire alignment again, starting from the Faraday.  We weren't that far off, so we didn't have to do anything too major.  Here's basically the procedure we used:

  • Using PZT 1 and 2 we directed the beam through the PRM aperture and through an aperture in front of PR2.  We also got good retro-reflection from PRM (with PRM free-hanging).  This completely determined our input pointing, and once it was done we DID NOT TOUCH the PZT mirrors any more.
  • The beam was fortunately still centered on PR2, so we didn't touch PR2.
  • Using PR3 we direct the beam through the BS aperture, through the ITMY aperture, and to the ETMY aperture.  This was accomplished by loosening PR3 and twisting it to adjust yaw, moving it forward/backwards to adjust the beam translation, and tapping the mirror mount to affect the hysteresis to adjust pitch.  Surprisingly this worked, and we were able to get the beam cleanly through the BS and Y arm apertures.  Reclamped PR3.
  • Adjusted ITMY biases (MEDM) to get Michelson Y arm retro-reflecting to BS.
  • Adjusting BS biases (MEDM) we directed the beam through the ITMX and ETMX apertures.
  • Adjusted ITMX biases (MEDM) to get Michelson X arm retro-reflecting to BS.

At this point things were looking good and we had Michelson fringes at AS.  Time to align SRC.  This is where things went awry yesterday.  Proceeded more carefully this time:

  • Loosened SR3 to adjust yaw pointing towards SRM.  We were pretty far off at SRM, but we could get mostly there with just a little bit of adjustment of SR3.  Got beam centered in yaw on SR2.
  • Loosened and adjusted SR2 to get beam centered in yaw on SRM.
  • Once we were centered on SR3, SR2, and SRM reclamped SR2/SR3.
  • Pitch adjustment was the same stupid stupid jabbing at SR2/3 to get the hysteresis to stick at an acceptable place.**
  • Looked at retro-reflection from SRM.  We were off in yaw.  We decided to adjust SRM pointing, rather than go through some painful SR2/3 iterative adjustment.  So unclamped SRM and adjusted him slightly in yaw to get the retro-reflection at BS.

At this point we felt good that we had the full IFO aligned.  We were then able to fairly quickly get the AS beam back out on the AS table.

We took at stab at getting the REFL beam situation figured out.  We confirmed that what we thought was REFL is indeed NOT REFL, although we're still not quite sure what we're seeing.  Since it was getting late we decided to close up and take a stab at it tomorrow, possibly after removing the access connector.

The main tasks for tomorrow:

  • Find ALL pick-off beams (POX, POY, POP) and get them out of the vacuum.  We'll use Jenne's new Suresh's old green laser pointer method to deal with POP.
  • Find all OPLEV beams and make sure they're all still centered on their optics and are coming out cleanly.
  • Center IPPOS and IPANG
  • Find REFL and get it cleanly out.
  • Do a full check of everything else to make sure there is no clipping and that everything is where we expect it to be.

Then we'll be ready to close.  I don't see us putting on heavy doors tomorrow, but we should be able to get everything mostly done so that we're ready on Monday.

** Comment: I continue to have no confidence that we're going to maintain good pointing with these crappy tip-tilt folding mirrors.

 

ELOG V3.1.3-